OpenStudy (anonymous):

Factor Completely: 3x2 − 21

OpenStudy (dumbcow):

factor out the 3 then use difference of squares

OpenStudy (anonymous):

\[x^2-y^2=(x+y)(x-y)\]

OpenStudy (anonymous):

3(x-7)(x+7)

OpenStudy (lalaly):

3(x^2 - 7) = 3(x-sqrt7)(x+sqrt7)

OpenStudy (anonymous):

i meant sqrt

OpenStudy (anonymous):

3(x-sqrt7)(x+sqrt7) this is what i meant

OpenStudy (dumbcow):

actually the answer could just be 3(x^2 -7) if they want all integers